OpenStudy (anonymous):

Really need lots of help on this :/ ill give out medals, best answer, and fan but please help im super behind on math and need to catch up quickly

OpenStudy (anonymous):

First, here is the existing map of current structures. It is important that the rollercoaster does not go through the foundation of any of these structures. Pick three points on the x–axis that the rollercoaster will dive underground or resurface. (Points that have a building on them are off-limits.) -1The rollercoaster will run from right to left, so work in that order when selecting points. 1st point:______ 2nd point:______ 3rd point:______

OpenStudy (anonymous):

-2) Using those points as zeros, construct the polynomial function, f(x), that will be the path of your rollercoaster. Show all of your work.

OpenStudy (anonymous):

-3) Using two of the theorems dicussed in the lesson, prove to the construction foreman that your function matches your graph. Use complete sentences. -4) Solve for the y–intercept for your function, f(x), and then graph your rollercoaster (you can use Geogebra). If your y–intercept is off the graph, give the coordinates of the y–intercept.

OpenStudy (anonymous):

@kewlgeek555

OpenStudy (johnweldon1993):

Well lets see....let say that we will go underground at x = 6....resurface at x = -2 ...and go underground again at x = -8

OpenStudy (johnweldon1993):

this means that our polynomial must have 0's at those points...so we have \[\large f(x) = (x - 6)(x + 2)(x + 8)\] *if you were to plug in the points we have...you would get your zeros here...so to find our polynomial function...multiply this out...can you do that?

OpenStudy (anonymous):

um yeah let me try

OpenStudy (anonymous):

f(x)=(6-6) (-2+2)(-8+8) is that good? im sorry im warning you i realy suck at this :/

OpenStudy (johnweldon1993):

Dont get down on yourself :) alright...so what I meant *which I should have been more clear on* is multiply out the function... \[\large f(x) = (x-6)(x+2)(x+8)\] have you used the foil method before? to multiply something like (a + b)(a +b) ??

OpenStudy (anonymous):

Nope :/

OpenStudy (anonymous):

Isnt it sorta like distributing? :s

OpenStudy (johnweldon1993):

Exactly! :) you take the first term in those first parenthesis...and multiply it by each term in the second parenthesis...then do that again for the second term in the first parenthesis.

OpenStudy (johnweldon1993):

so we'll start with this... we have \[\large f(x) = (x - 6)(x + 2)(x +8)\] Lets break it up \[\large (x - 6)(x + 2)\] can you distribute (foil) this?

OpenStudy (anonymous):

Yay ok awesome haha ok give me a sec

OpenStudy (anonymous):

would it be x^2+2x-6x-12 ?

OpenStudy (anonymous):

Are you in Connections academy? Is this Algebra 2 Foundations b?

OpenStudy (anonymous):

Yes c: and im in Algebra 2 idk what Foundations b is :))

OpenStudy (johnweldon1993):

@omgkelley hey I'm back...and yes you distributed correctly....just look though...there is a 2x and a -6x....what should we do?

OpenStudy (anonymous):

can you help me with Unit three lesson 3?

OpenStudy (anonymous):

umm make them smaller? so would it be x & -3 ? @johnweldon1993

OpenStudy (usarmy3947):

do u still need help?

OpenStudy (anonymous):

Yessss :c

OpenStudy (usarmy3947):

ok lets see and also did i help in the youtube favor?

OpenStudy (anonymous):

I've been stuck ont his for over a week :/

OpenStudy (anonymous):

Please help guys :/ If i dont finish this ill most probably fail Math

OpenStudy (anonymous):

what @johnweldon1993 was doing is a good start. Let's continue with what he asked you to do with the 2x-6x So \[2x-6x=x(2-6)=?\]

OpenStudy (anonymous):

umm okay im lost

OpenStudy (anonymous):

^_^ no worries \[2x-6x=-4x\]

OpenStudy (anonymous):

All I did above was I pulled the x out in front to make the arithmetic more visible \[x(2-6)=x(-4)=-4x\] Can you see where I got that?

OpenStudy (anonymous):

um noo i saw that you solved the 2x-6x but you got me lost up there ^

OpenStudy (anonymous):

And this was all from the first part of the polynomial he created up top \[f(x)=(x-6)(x+2)(x+8)\] We're just looking at the \[(x-6)(x+2)\] part first. \[(x-6)(x+2)=x^2-6x+2x-12 \\ \hspace{2.4cm} =x^2-4x-12\] So \[f(x)=\big((x-6)(x+2)\big)(x+8) \\ \ \\ \qquad = (x^2-6x+2x-12)(x+8)\]

OpenStudy (anonymous):

But you understand the \[2x-6x=-4x\] if I don't do the "pull x out" thing?

OpenStudy (anonymous):

Yes i understand that part :)

OpenStudy (anonymous):

then bueno! ^_^ looking at the next part of the distributing then, you have \[f(x)=\big((x-6)(x+2)\big)(x+8) \\ \ \\ \qquad = (x^2-4x-12)(x+8)\] So you just do like you did before, multiplying each of the terms in the quadratic by x, and then again by 8, then simplifying it like did before

OpenStudy (anonymous):

oh okay! easy :D

OpenStudy (anonymous):

Hey, I'm really sorry, but I have to run to an appointment. @SolomonZelman Can you help out with the rest of this problem?

OpenStudy (anonymous):

oh no D:

OpenStudy (anonymous):

In case he's busy, for the last part of the problem you just plug in x=0 into the polynomial you get from simplifying to find the y intercept, so every term with an x in it disappears and you're just left with a number ^_^ Good luck!! You'll be awesome!!!!!! I'll be back a bit later if you still need help with this ^_^

OpenStudy (anonymous):

I probably wiil haha i'll wait up C: and thanks for being so patient with me haha

OpenStudy (anonymous):

Anytime! ^_^

OpenStudy (anonymous):

@tester97 have you done this yet?

OpenStudy (tester97):

what exactly is the question?

OpenStudy (anonymous):

Theres originally 5 but im doing the alternate assessment which is just 4

OpenStudy (tester97):

oh sorry i didnt do this

undeadknight26 (undeadknight26):

Solo is off line..so i will try what i the question?

OpenStudy (anonymous):

Okay hold on

undeadknight26 (undeadknight26):

Lol there are some confusing things going on in this question...i can't find the question...

OpenStudy (anonymous):

The rollercoaster will run from right to left, so work in that order when selecting points. 1st point:__6____ 2nd point:___-2___ 3rd point:____-8__ Using those points as zeros, construct the polynomial function, f(x), that will be the path of your rollercoaster. Show all of your work.

undeadknight26 (undeadknight26):

ok so a function is f(x) = mx + b where m is your slope and b is your y.

OpenStudy (anonymous):

LLol yeah ik, sorry ^ those are the first two questions. there are four in total

undeadknight26 (undeadknight26):

Lets try and work them out piece by piece i think i know how to do this :3

undeadknight26 (undeadknight26):

So what do you think ur function is?

OpenStudy (anonymous):

omg okay yay! youre such a life saver . umm i dont know im lost on this whole lesson

undeadknight26 (undeadknight26):

Lol i said i think :D lets just remember that :D

OpenStudy (anonymous):

Still no one elsee seems to even understand the question haha

undeadknight26 (undeadknight26):

lol i don't understand why not...all u need to do is make a function...yet finding the slope...i forgot how to find that...do you remember?

OpenStudy (anonymous):

Ummm isnt it y2-y1 divided by x2-x1 ?

undeadknight26 (undeadknight26):

oh yes...but ur points are:(1,6)(2,-2)(3,-8)? I believe that is it right?

undeadknight26 (undeadknight26):

if so...i has no idea this is quite confusing now lol...Hmmm...

undeadknight26 (undeadknight26):

@BangkokGarrett this is very hard can u help me with this?

OpenStudy (anonymous):

Well if you have any other points ill take them c:

undeadknight26 (undeadknight26):

Maffs told u the answer yet it is like deciphering a futuristic code...

undeadknight26 (undeadknight26):

What grade are you in?

OpenStudy (anonymous):

Hahaha this assignment is sooo confusing

OpenStudy (anonymous):

10th but im taking 11th grade classes

undeadknight26 (undeadknight26):

lol im in 9th...

OpenStudy (anonymous):

Hey @omgkelley Sorry that our paths didn't cross again to finish the problem and that you had to close the question :/ What @undeadknight26 said about the slope isn't what this question is looking for, because it's a cubic polynomial. Instead of a straight line, it'll look something like this.  @johnweldon1996 helped choose the points well, so that your roller coaster would avoid all of the buildings, and not resurface in any foundation (like the problem wanted). Looking at the polynomial that you came up with again, \[f(x)=(x-6)(x+2)(x+8) \\ \ \ \ =(x^2-4x-12)(x+8) \\ \ \ \ =(x)(x^2-4x-12)+(8)(x^2-4x-12)\] (you can finish simplyfying ^_^ Just distribute the x in the first expression, the 8 in the second expression, and add up all the like terms) you can see that the roots of the equation (where the line of the graph crosses the x axis - aka, the points where the y axis is zero) were developed by defining the ground entry points, (8,0) and (-6,0), and exit point (-2,0) as the roots of a cubic polynomial (a polynomial where the highest power x is raised to is 3 - when you finish simplifying the above term, you'll see that there's an x^3 term). Since you know (by definition) that where the function \[f(x)=(x-6)(x+2)(x+8)\] exists at y=0, those will be your roots (shown below). If it helps you visualize it better, you can write the function in terms of y and x. \[y=(x-6)(x+2)(x+8)\] It's the same thing. Now, again knowing that where the roller coaster enters and exits the ground all exist at y=0, all you have to do is set y=0 to prove these are the points you need for your function! Since the ground is at \[y=0\] then \[(x-6)(x+2)(x+8)=0\] giving the three distinct "roots," or values of x that satisfy the equation (chosen from points on the graph initially). Those are, as chosen, \[\qquad \begin{array}{l} x-6=0 & \longrightarrow & x=6 & \rightarrow & \text{corresponds to point (6,0)} \\x+2=0 & \longrightarrow & x=-2 & \rightarrow & \text{corresponds to point (-2,0)} \\ x+8 = 0 & \longrightarrow & x=-8 & \rightarrow & \text{corresponds to point (-8,0)} \end{array}\] (To convince yourself, you can plug in each of those values of x one at a time and make sure the function equals zero for those values. ie \[f(-8) = (-8-6)(-8+2)(-8+8)=(-14)(-6)(0)=0 \ \large \color{green}\checkmark\] The last part of the question that asks you to find the y-intercept (by definition, where the line crosses the y axis; and also by definition where the x-axis equals zero) can be completed in a similar way to plugging in one of the values of the roots. All you have to do is substitute in x=0 for all x 's in the equation. \[f(0) = (0-6)(0+2)(0+8)=(-6)(2)(8)= \text{y-intercept}\] This is exactly like finding the y intercept of a normal straight line. For instance, if you had a straight line, with the equation of a line \[y(x)=mx+b\] you know that b is the y-intercept, because at x=0 \[ y(0)=m(0)+b=b=\text{y-intercept}\] In your equation there are just a few more x terms. Hope that this helped - holler if you have any questions! And good luck!! ^_^